Lura:

Math help plz -5(2x-5)-3x-5=7

Lura:

bouta say "dude what da f-"

XLynx:

-10x+25-3x-5 = 7 -13x+20 = 7 -13x = -13 x = 1

Extrinix:

Forgot to explain, xd

XLynx:

i love math lol

XLynx:

i have 790 SAT math

Extrinix:

But here, Simplify Add similar values together Subtract both sides by 20 Divide by -13 on both sides

Lura:

Thanks guys ^^

Extrinix:

Lol thank Lyn, not me (I just made sure he didn’t get warned 😉)

Lura:

but i want to ;w;

XLynx:

thanks preki :)